OpenStudy (anonymous):

2(x + 6) − 10 = 12 + 4(x − 1)

OpenStudy (mayankdevnani):

2x+12-10=12+4x-4 2x+2=8+4x -2x=6 -x=6/2=3 x=-3

OpenStudy (anonymous):

i agrre

OpenStudy (anonymous):

i meant agree

OpenStudy (mayankdevnani):

thnx